# AutoP2C: An LLM-Based Agent Framework for Code Repository Generation from Multimodal Content in Academic Papers

### v1

@@ 4.2. Benchmark and Baseline
The P2C task represents a novel challenge that extends beyond traditional code generation by requiring the integration of multimodal content into coherent, multi-file code repositories. Given the absence of specialized benchmarks or frameworks designed specifically for this task, we first establish a new benchmark comprising eight recent research papers all published after 2024 (Fontana et al., 2024; Nasiri and Garraghan, 2025; Nazari et al., 2024; Xu et al., 2024b; Liang et al., 2024; Doshi and Kim, 2024; Azimi et al., 2024; Linse et al., 2023). These papers are selected from paperswithcode.com to evaluate the capability of AutoP2C in processing diverse modal datasets across various ML domains. The benchmark spans six ML tasks (including training strategy optimization (Xu et al., 2024b), node classification (Nazari et al., 2024; Liang et al., 2024), model compression (Doshi and Kim, 2024), parameter-efficient fine-tuning (Azimi et al., 2024; Nasiri and Garraghan, 2025), network pruning (Fontana et al., 2024), and image classification (Linse et al., 2023)) and three modal types of datasets (containing computer vision (Fontana et al., 2024; Xu et al., 2024b; Doshi and Kim, 2024; Linse et al., 2023), natural language processing (Nasiri and Garraghan, 2025; Azimi et al., 2024), and graph (Nazari et al., 2024; Liang et al., 2024) related datasets). We also established baselines using two state-of-the-art LLMs: OpenAI’s recently released o1 (Zhong et al., 2024) and DeepSeek-R1 (Guo et al., 2025), both known for their advanced code generation capabilities.
@@ 4.4.1. Evaluation for Functional Correctness
The results in Table 1 demonstrate the advantages of AutoP2C in both implementation feasibility and functional fidelity compared to the baselines. While OpenAI-o1 and DeepSeek-R1 successfully generate executable code for only one paper (“No More Adam: …” (Xu et al., 2024b) ), AutoP2C produces functional implementations across all eight research papers in our benchmark. This contrast in success rate (100% versus 12.5%) underscores the effectiveness of AutoP2C in performing the P2C task. Even in the single case where the baseline models produce executable code, AutoP2C achieves superior performance. Specifically, for “No More Adam: …” (Xu et al., 2024b), our AutoP2C delivers 91.4% accuracy compared to 81.8% for o1 (9.6% improvement) and 75.3% for R1 (16.1% improvement). The relative performance metric further validates the effectiveness of our framework, with AutoP2C achieving between 89.8% and 122.0% of the performance of the original implementations in all eight papers. Notably, in two cases, the performance of the code repositories generated by AutoP2C exceeds that of the code of the original papers, namely “Entropy Aware..” (Nazari et al., 2024) (103.0%) and “Convolutional…” (Linse et al., 2023) (122.0%). The consistent high relative performance (averaging 99.5% across all papers) confirms that AutoP2C’s staged generation paradigm ensures both compliance with paper specifications and operational integrity in real-world execution. Additionally, we provide a case study of AutoP2C in Fig. 3, demonstrating how our approach successfully transforms multimodal research content into executable code through its four-stage pipeline.

### v2

@@ 3.6. Iterative Feedback-Driven Implementation
Table 2. Average Replication Scores (in %) of different baselines on PaperBench Code-Dev benchmark using o3-mini-high. The results of the baseline in the first three lines are all from the paper (Seo et al., 2025). ‘$\pm$’ represents the standard error of the mean.
Baseline Replication Score (%)
BasicAgent $5.1\pm 0.8$
IterativeAgent $16.4\pm 1.4$
PaperCoder (Seo et al., 2025) $44.2$
AutoP2C $49.2\pm 14.0$
@@ 4.2. Benchmark and Baseline
The P2C task represents a novel challenge that extends beyond traditional code generation by requiring the integration of multimodal content into coherent, multi-file code repositories. To provide a comprehensive evaluation, we consider two complementary benchmarks, Paper2Repo, and PaperBench Code-Dev (Starace et al., 2025), each emphasizing a different aspect of repository quality. (1) To evaluate how faithfully AutoP2C implements diverse models and algorithmic designs described in each paper, we constructed a brand-new benchmark called Paper2Repo, which comprises eight recent research papers all published after 2024 (Fontana et al., 2024; Nasiri and Garraghan, 2025; Nazari et al., 2024; Xu et al., 2024b; Liang et al., 2024; Doshi and Kim, 2024; Azimi et al., 2024; Linse et al., 2023). Each paper contains all the content required for the reproduction of the code repository. These papers are selected from paperswithcode.com, and each has a corresponding GitHub repository as the upper bound for comparison. The benchmark spans six ML tasks (including training strategy optimization (Xu et al., 2024b), node classification (Nazari et al., 2024; Liang et al., 2024), model compression (Doshi and Kim, 2024), parameter-efficient fine-tuning (Azimi et al., 2024; Nasiri and Garraghan, 2025), network pruning (Fontana et al., 2024), and image classification (Linse et al., 2023)) and three modal types of datasets (containing computer vision (Fontana et al., 2024; Xu et al., 2024b; Doshi and Kim, 2024; Linse et al., 2023), natural language processing (Nasiri and Garraghan, 2025; Azimi et al., 2024), and graph (Nazari et al., 2024; Liang et al., 2024) related datasets). We also established baselines using two state-of-the-art LLMs: OpenAI’s recently released o1 (Zhong et al., 2024) and DeepSeek-R1 (Guo et al., 2025), both known for their advanced code generation capabilities. (2) To further assess the completeness of the code repositories generated by AutoP2C, we evaluated our framework on OpenAI’s newly released PaperBench Code-Dev benchmark (Starace et al., 2025). We compared AutoP2C with several established baselines introduced in (Starace et al., 2025), as well as PaperCoder, a recent multi-agent framework expressly designed to create code repositories from research papers.
@@ 4.4.1. Evaluation for Functional Correctness
The results in Table 1 demonstrate the advantages of AutoP2C in both implementation feasibility and functional fidelity compared to the baselines. While OpenAI-o1 and DeepSeek-R1 successfully generate executable code for only one paper (“No More Adam: …” (Xu et al., 2024b) ), AutoP2C produces functional implementations across all eight research papers in our benchmark. This contrast in success rate (100% versus 12.5%) underscores the effectiveness of AutoP2C in performing the P2C task. Even in the single case where the baseline models produce executable code, AutoP2C achieves superior performance. Specifically, for “No More Adam: …” (Xu et al., 2024b), our AutoP2C delivers 91.4% accuracy compared to 81.8% for o1 (9.6% improvement) and 75.3% for R1 (16.1% improvement). The relative performance metric further validates the effectiveness of our framework, with AutoP2C achieving between 89.8% and 122.0% of the performance of the original implementations in all eight papers. Notably, in two cases, the performance of the code repositories generated by AutoP2C exceeds that of the code of the original papers, namely “Entropy Aware..” (Nazari et al., 2024) (103.0%) and “Convolutional…” (Linse et al., 2023) (122.0%). The consistently high relative performance (averaging 99.5% across all papers) confirms that AutoP2C’s staged generation paradigm ensures both compliance with paper specifications and operational integrity in real-world execution. Additionally, we provide a case study of AutoP2C in Fig. 3, demonstrating how our approach successfully transforms multimodal research content into executable code through its four-stage pipeline.
@@ 4.4.2. Evaluation for Structural Completeness
4.4.3. Evaluation for PaperBench Code-Dev
To gauge how well AutoP2C generalizes beyond our own curated corpus, we further benchmark it on PaperBench Code-Dev, a public suite that automatically grades end-to-end reproduction of research code. As summarised in Table  2, AutoP2C attains the highest average replication score, $49.2\pm 14.0\%$, surpassing PaperCoder (44.2%) by roughly five percentage points and outperforming the lightweight BasicAgent ($5.1\pm 0.8\%$) and IterativeAgent ($16.4\pm 1.4\%$) by wide margins. These results confirm that our structured multi-agent pipeline translates more of a paper’s specification into functional, runnable code than existing single- or few-agent baselines.